# PII-Bench: Evaluating Query-Aware Privacy Protection Systems

### v1

@@ 4.5 In-depth Performance Analysis
Impact of Model Scale. Analysis reveals consistent performance gaps between large and small models across tasks. Small-scale variants show 13-33% lower F1 scores in PII detection, with wider gaps in the query-related task (Llama3.1-SLM: 0.328 F1 vs. GPT4o: 0.627 F1) and query-unrelated masking task (0.42-0.54 F1 vs. 0.70-0.72 F1).
PII Type Performance. As shown in Fig. 5, models perform better at recognizing structured information (PER, CODE) compared to contextual entities (LOC, ORG), suggesting models have stronger capability in identifying patterns with clear structural characteristics.
Factors Influencing Performance. Fig. 4 reveals several critical factors affecting model accuracy: performance degrades sharply beyond 5 subjects (F1 drops from 0.85 to 0.52), 33 PII entities, or 3000 characters in text length. Query-related entity count shows modest impact, with gradual decline from 1 to 7 entities.
@@ 5 Conclusion
This paper introduces PII-Bench, a comprehensive evaluation framework comprising 2,842 test samples, along with a query-unrelated PII masking strategy. Our evaluation reveals that while current LLMs achieve strong performance in basic PII detection (F1>0.90), they show limited capability in query-relevance assessment (F1<0.63) and struggle with complex multi-subject scenarios. Small-scale models demonstrate substantially lower performance across all tasks. These findings establish important benchmarks for privacy-preserving systems and highlight critical challenges in intelligent PII handling.
@@ A.2 PII Types
Unlike the PII types presented by Papadopoulou et al.’s (2022), our classification does not include the MISC category. This exclusion is due to the ambiguous definition of the MISC category and its unclear boundaries with other categories.
The definitions of the seven categories are as follows:
@@ B.3 Entity Categories and Generation Methods
Appendix C Human Evaluation Details
Appendix D Experiments Details
D.1 Evaluation Metrics
$P_{strict}(p_{i},g_{j})=\frac{|E_{p_{i}}\cap E_{g_{j}}|}{|E_{p_{i}}|}$ (1)
$R_{strict}(p_{i},g_{j})=\frac{|E_{p_{i}}\cap E_{g_{j}}|}{|E_{g_{j}}|}$ (2)
where $E_{p_{i}}$ and $E_{g_{j}}$ are the sets of entity-type pairs.
2. Entity-only Matching: Only entity spans need to match:
$P_{ent}(p_{i},g_{j})=\frac{|S_{p_{i}}\cap S_{g_{j}}|}{|S_{p_{i}}|}$ (4)
$R_{ent}(p_{i},g_{j})=\frac{|S_{p_{i}}\cap S_{g_{j}}|}{|S_{g_{j}}|}$ (5)
where $S_{p_{i}}$ and $S_{g_{j}}$ are the sets of entity spans.
$P_{span}$, $R_{span}$, and $F1_{span}$ are computed analogously.
$P_{query}=\frac{|\mathcal{E}_{p}\cap\mathcal{E}_{g}|}{|\mathcal{E}_{p}|}$ (11)
$R_{query}=\frac{|\mathcal{E}_{p}\cap\mathcal{E}_{g}|}{|\mathcal{E}_{g}|}$ (12)
$F1_{query}=\frac{2\cdot P_{query}\cdot R_{query}}{P_{query}+R_{query}}$ (13)
$F1_{fuzzy}=\frac{2\cdot P_{fuzzy}\cdot R_{fuzzy}}{P_{fuzzy}+R_{fuzzy}}$ (16)
D.2 Additional Results
GPT4o Llama3.1 Qwen2.5 Llama3.1-SLM Qwen2.5-SLM
Method F1 RougeL-F F1 RougeL-F F1 RougeL-F F1 RougeL-F F1 RougeL-F
Basic Method
Naive 0.600 0.602 0.611 0.614 0.596 0.603 0.240 0.333 0.405 0.413
Advanced Method
Self-CoT 0.675 0.681 0.638 0.643 0.626 0.632 0.354 0.362 0.392 0.397
PS-CoT 0.618 0.620 0.624 0.631 0.636 0.643 0.291 0.300 0.431 0.436
w/ Extra Information
D.3 Prompt Details
Appendix E PII Annotation System
E.1 PII Detection Interface
E.2 Query-Related Detection Interface
E.3 Query-unrelated PII Masking Visualization
E.4 Annotation Guidelines and Quality Control
To ensure annotation consistency and quality, we established comprehensive guidelines and implemented several control measures:
  • Entity Span Guidelines: Annotators must select the minimal text span that completely captures the PII entity while maintaining semantic integrity.
  • Inter-annotator Agreement: Each sample is independently annotated by multiple annotators, with disagreements resolved through majority voting or expert review.
  • Validation Checks: The platform implements automatic validation rules to detect potential inconsistencies or missing annotations.
  • Iterative Refinement: Regular review sessions are conducted to discuss challenging cases and update guidelines based on annotator feedback.
For quality assurance, we randomly sampled 10% of the annotations for expert review, achieving an inter-annotator agreement of 95.1% for PII detection and 91.5% for query-related detection across all annotators.
@@ E.4 Annotation Guidelines and Quality Control
CODE Social Security Number Rule-based XXX-XX-XXXX
@@
DATETIME Date Rule-based YYYY-MM-DD
Time Rule-based HH:MM:SS

### v2

@@ 4.7 In-depth Performance Analysis
Factors Influencing Performance. Figure 4 reveals several critical factors affecting model accuracy: performance degrades sharply beyond 5 subjects (F1 drops from 0.85 to 0.52), 33 PII entities, or 3000 characters in text length. Query-Related entity count shows modest impact, with gradual decline from 1 to 7 entities.
@@ 5 Conclusion
This paper introduces PII-Bench, a comprehensive evaluation framework comprising 2,842 test samples across 7 PII types with 55 fine-grained subcategories, and proposes a query-unrelated PII masking strategy to balance privacy protection with LLM utility. Our empirical evaluation reveals that while advanced LLMs demonstrate strong performance in basic PII detection, they exhibit substantial limitations in query-relevance assessment. Small-scale models show considerably larger performance gaps across all evaluation tasks. These findings establish foundational benchmarks and exposes critical challenges in privacy-aware PII handling.
@@ A.2 PII Types
Unlike Papadopoulou et al.’s (2022) classification, we exclude the MISC category due to its ambiguous definition and unclear boundaries. Our taxonomy comprises seven categories:
@@ B.3 Entity Categories and Generation Methods
CODE Social Security Number Rule-based XXX-$X$-XXXX
@@
DATETIME Date Rule-based YYYY-$M$-$D$
Time Rule-based $H$:$M$:$S$
@@
Appendix C Consistency Optimization Details
The consistency optimization process is critical for ensuring that randomly sampled PII entities form logically coherent and realistic profiles. We employ GPT-4-0806 with carefully designed prompts to identify and resolve conflicts while preserving the diversity and coverage of PII types.
C.1 Single-Subject Consistency Optimization
For single-subject descriptions, the optimization process addresses intra-subject conflicts arising from incompatible entity combinations. Common conflict patterns include:
  • Temporal inconsistencies: Age incompatible with work experience duration, education level, or career stage (e.g., a 23-year-old with 15 years of work experience).
  • Professional mismatches: Occupation inconsistent with education background or salary range (e.g., a high school graduate working as a licensed physician).
  • Geographic contradictions: Work location distant from residential address without supporting evidence (e.g., daily commute spanning different continents).
  • Financial implausibility: Income, expenses, and savings that violate basic economic constraints (e.g., monthly expenses exceeding monthly income by orders of magnitude).
The optimization prompt (Figure 10) instructs the model to: (1) identify logical conflicts among sampled entities; (2) modify only the entity values while preserving their PII type and category labels; (3) maximize entity retention to maintain dataset richness. For instance, given a conflict between “Age: 22 years” and “Work Experience: 15 years as Senior Engineer”, the optimization adjusts the work experience to “2 years as Junior Engineer” rather than removing the entity entirely, thereby preserving both the PII type (DEM) and the entity category (Work Experience).
C.2 Multi-Subject Consistency Optimization
Multi-subject optimization extends the single-subject process by introducing relationship-aware constraints. Consistency rules are enforced based on three relationship categories:
  • Intersection relationships (e.g., colleagues, classmates): Subjects must share critical attributes such as organization name, work location for colleagues, or educational institution for classmates. The optimization verifies attribute alignment and adjusts inconsistent entities accordingly.
  • Hierarchical relationships (e.g., parent-child, supervisor-subordinate): Structural constraints are enforced such as age differences (parent at least 18-20 years older than child), authority levels (supervisor holding higher position than subordinate), and derived attributes (children inheriting nationality or ethnicity from parents when culturally appropriate).
  • Non-intersection relationships: Subjects maintain independent entity sets with no enforced dependencies, though internal consistency within each subject’s profile is still verified.
The multi-subject prompt (Figure 11) provides relationship context and requires the model to output separate, coordinated entity sets for each subject. This relationship-aware approach ensures that multi-subject descriptions reflect realistic interpersonal dynamics and maintain narrative coherence across profiles.
Appendix D Query Scenario Design Details
The scenario design phase bridges the gap between abstract entity selection and natural user queries by grounding queries in realistic application contexts. This process is essential for evaluating whether privacy protection systems can identify query-relevant PII in authentic usage scenarios.
D.1 Domain-Specific Scenario Generation
Given a set of selected entities $\mathcal{E}_{q}$, we employ a structured prompting strategy to generate diverse scenario contexts. The prompt instructs the LLM to:
  1. 1.
    Analyze entity semantics: Determine the thematic domain implied by the selected entities (e.g., education and work experience suggest career-related scenarios; medical conditions and medications suggest healthcare scenarios).
  2. 2.
    Generate scenario candidates: Propose multiple scenario types that naturally require all selected entities, ensuring thematic diversity across the dataset. Example scenarios include career planning, medical consultation, financial advisory, legal counseling, academic mentoring, and housing applications.
  3. 3.
    Ensure entity necessity: Verify that each selected entity is meaningfully incorporated into the scenario rather than superficially included, so that query-relevant PII labels reflect genuine information dependencies.
The scenario generation prompt is provided in Figure 12. By requiring coverage of all selected entities while maintaining scenario diversity, this approach produces queries that realistically reflect specialized user information needs across domains.
D.2 Scenario Validation and Filtering
Generated scenarios undergo validation to ensure quality and diversity:
  • Completeness check: Verify that all entities in $\mathcal{E}_{q}$ are semantically incorporated into the scenario context.
  • Realism assessment: Confirm that the scenario represents plausible real-world interactions with AI systems.
  • Diversity enforcement: Track scenario distribution and reject over-represented scenario types to maintain balanced domain coverage.
This rigorous scenario design process ensures that PII-Bench queries reflect authentic information needs across diverse domains, providing a realistic testbed for evaluating query-aware privacy protection systems.
Appendix E Human Evaluation Details
Appendix F Experiment Details
F.1 Evaluation Metrics
$P_{strict}(p_{i},g_{j})=\frac{|E_{p_{i}}\cap E_{g_{j}}|}{|E_{p_{i}}|}$ (1)
$R_{strict}(p_{i},g_{j})=\frac{|E_{p_{i}}\cap E_{g_{j}}|}{|E_{g_{j}}|}$ (2)
where $E_{p_{i}}$ and $E_{g_{j}}$ are the sets of entity-type pairs.
2. Entity-only Matching: Only entity spans need to match:
$P_{ent}(p_{i},g_{j})=\frac{|S_{p_{i}}\cap S_{g_{j}}|}{|S_{p_{i}}|}$ (4)
$R_{ent}(p_{i},g_{j})=\frac{|S_{p_{i}}\cap S_{g_{j}}|}{|S_{g_{j}}|}$ (5)
where $S_{p_{i}}$ and $S_{g_{j}}$ are the sets of entity spans.
$P_{span}$, $R_{span}$, and $F1_{span}$ are computed analogously.
$P_{query}=\frac{|\mathcal{E}_{p}\cap\mathcal{E}_{g}|}{|\mathcal{E}_{p}|}$ (11)
$R_{query}=\frac{|\mathcal{E}_{p}\cap\mathcal{E}_{g}|}{|\mathcal{E}_{g}|}$ (12)
$F1_{query}=\frac{2\cdot P_{query}\cdot R_{query}}{P_{query}+R_{query}}$ (13)
$F1_{fuzzy}=\frac{2\cdot P_{fuzzy}\cdot R_{fuzzy}}{P_{fuzzy}+R_{fuzzy}}$ (16)
F.2 Privacy-Utility Tradeoff Metrics
We develop a comprehensive evaluation framework to quantify the tradeoff between privacy protection and query utility across different PII masking strategies.
F.2.1 Privacy Protection Metric
Let $\mathcal{E}=\{e_{1},...,e_{n}\}$ denote the set of PII entities in the original text $T_{o}$, and $T_{m}$ represent the masked text. The privacy score $P$ measures the proportion of PII entities successfully protected:
$P=1-\frac{\sum_{e\in\mathcal{E}}C(e,T_{m})}{\sum_{e\in\mathcal{E}}C(e,T_{o})}$ (17)
where $C(e,T)$ counts occurrences of entity $e$ in text $T$. A score of $P=1$ indicates complete protection, while $P=0$ indicates no protection.
F.2.2 Utility Preservation Metrics
To measure utility preservation, we employ two complementary approaches:
Semantic Similarity
We compute embedding-based similarity between responses generated from masked prompts ($R_{m}$) and unmasked prompts ($R_{o}$):
$U_{s}=\cos(\mathbf{v}_{R_{m}},\mathbf{v}_{R_{o}})$ (18)
where $\mathbf{v}_{R}$ is the text embedding generated by BGE-M3 Multi-Granularity (2024), and $\cos(\cdot,\cdot)$ computes cosine similarity. This metric captures semantic preservation independent of exact wording.
LLM-as-Judge Evaluation
We employ Claude-3.7-Sonnet to assess response quality through direct comparison:
$U_{l}=Judge(R_{m}^{1},R_{m}^{2},R_{o})$ (19)
where $R_{m}^{1}$ and $R_{m}^{2}$ represent responses from two different masking strategies, and $R_{o}$ is the reference response from unmasked text. The judge assigns numerical ratings $r\in[1,10]$ to each masked response based on how well it preserves the query intent compared to the reference. To mitigate position bias, we randomly alternate the presentation order of responses.
F.2.3 Balanced Metric
To quantify the overall effectiveness of masking strategies, we compute a balanced score $B$ that combines privacy protection and utility preservation:
$B=\alpha P+(1-\alpha)U$ (20)
where $\alpha\in[0,1]$ is a weighting parameter that determines the relative importance of privacy versus utility. In our experiments, we use $\alpha=0.5$ to assign equal importance to both aspects.
The complete prompt template for LLM-as-Judge evaluation is provided in Figure 20.
F.3 Additional Experiments with Smaller Language Models
We have conducted a comprehensive evaluation on smaller, deployment-ready models (0.5B-3B parameters) to assess their viability for on-device PII protection.
PII-Single PII-Multi PII-Hard PII-Distract
Qwen2.5-0.5B 0.002 0.0029 0.002 0.0013 0.0042 0.0013 0.0034 0.0042 0.0034 0.009 0.0207 0.009
Qwen2.5-1.5B 0.1846 0.2069 0.1865 0.1057 0.1794 0.1071 0.1474 0.1976 0.1502 0.0728 0.1549 0.0733
Qwen2.5-3B 0.5929 0.6289 0.5962 0.6189 0.7067 0.6212 0.5202 0.5921 0.5237 0.3717 0.6293 0.3731
Table 11: Performance of baseline models under the PII Detection task.
Qwen2.5-0.5B Qwen2.5-1.5B Qwen2.5-3B
Method F1 RougeL-F F1 RougeL-F F1 RougeL-F
Basic Method
Naive 0.0041 0.0041 0.2131 0.2185 0.1691 0.1699
Advanced Method
Self-CoT 0.009 0.0095 0.2002 0.2051 0.2884 0.2909
Auto-CoT(3-shot) 0.0258 0.0268 0.2383 0.2466 0.381 0.384
Self-Consistency 0.0018 0.0018 0.1057 0.1088 0.2694 0.2774
PS-CoT 0.0088 0.009 0.1887 0.1959 0.293 0.2956
w/ PII Detection
Naive w/ Choice 0.3978 0.3986 0.4357 0.4357 0.542 0.5437
Qwen2.5-0.5B Qwen2.5-1.5B Qwen2.5-3B
Method F1 RougeL-F F1 RougeL-F F1 RougeL-F
Basic Method w/ PII Detection
Naive 0.0016 0.0025 0.0912 0.1018 0.3257 0.3321
Advanced Method w/ PII Detection
Self-CoT 0.0016 0.0029 0.0846 0.0945 0.3764 0.3841
Auto-CoT(3-shot) 0.0019 0.0032 0.0927 0.1038 0.4247 0.4327
Self-Consistency 0.0016 0.0029 0.0907 0.1020 0.3414 0.3486
PS-CoT 0.0016 0.0025 0.0920 0.1033 0.3765 0.3840
w/ PII Detection
Naive w/ Choice 0.0011 0.0017 0.0722 0.0800 0.3918 0.3994
Table 13: Performance comparison on the Query-Unrelated PII Masking task (PII-single and PII-multi datasets).
As shown in Tables 1213, and 11, our results reveal a significant performance gap between deployment-ready models and larger proprietary systems. The 0.5B and 1.5B models exhibit extremely limited capabilities across all tasks (with F1 scores generally below 0.2), rendering them practically unusable for real-world PII protection. While the 3B model shows modest potential (F1 scores approaching 0.6 on simpler datasets), its performance remains substantially inferior to larger models, particularly on challenging scenarios (e.g., PII-Hard, PII-Distract). These findings underscore the tension between privacy goals and model capabilities: while smaller on-device models would be preferable from a privacy perspective, they currently lack the sophistication needed for reliable PII management.
F.4 Supplementary Results on PII-Real Dataset
To empirically validate PII-Bench’s ability to capture real-world privacy challenges, we constructed PII-Real, a validation dataset based on authentic biographical information.
F.4.1 Dataset Construction
We selected 20 prominent AI researchers from the AMiner AI2000 ranking [^1] https://www.aminer.cn/ai2000 and manually curated their professional profiles from publicly available sources including conference websites, institutional pages, and academic publications. Each profile underwent expert annotation by the authors following our established PII taxonomy, ensuring consistency with PII-Bench guidelines.
Unlike PII-Bench’s automated query generation, PII-Real features human-written queries crafted by domain experts to reflect authentic information needs. We designed queries spanning career planning, research collaboration, academic advising, and startup leadership scenarios for each profile. The resulting dataset comprises 100 instances with manually annotated PII entities, providing an authentic testbed for validating synthetic data quality.
Tables 1718, and 19 present experimental results across three evaluation tasks.
F.4.2 Consistency Analysis
We quantified alignment between PII-Single and PII-Real using Spearman’s rank correlation coefficient ($\rho$) and Performance Consistency Score (PCS).
Performance Consistency Score
PCS measures absolute performance alignment between two datasets. For a given model and evaluation metric, PCS is defined as:
$PCS=1-\frac{|F1_{single}-F1_{real}|}{max(F1_{single},F1_{real})}$ (21)
where $F1_{single}$ represents the F1 score on PII-Single dataset and $F1_{real}$ represents the F1 score on PII-Real dataset. A PCS value of 1 indicates perfect performance consistency, while lower values indicate greater deviation between synthetic and real-world scenarios.
Ranking Consistency
Table 14 presents Spearman’s $\rho$ values across all evaluation tasks. The exceptionally high correlation coefficients ($\rho>0.988$, $p<0.001$) demonstrate that model rankings remain consistent between synthetic and real-world datasets, indicating PII-Bench reliably predicts relative model performance in authentic privacy scenarios.
Task Spearman’s $\rho$ P-value
PII Detection 0.9880 < 0.001
Query-Related Detection 0.9982 < 0.001
Query-Unrelated Masking 0.9982 < 0.001
Table 14: Spearman’s rank correlation between PII-Single and PII-Real datasets across evaluation tasks.
Performance Consistency
Tables 15 and 16 present PCS values quantifying absolute performance alignment. For PII detection (Table 15), all models achieve PCS values exceeding 0.967, with an overall average of 0.980. For query-related detection and masking tasks (Table 16), PCS values consistently exceed 0.947 across all model-strategy combinations, averaging 0.966. These high consistency scores confirm that synthetic single-entity scenarios accurately capture the challenges present in real-world privacy protection tasks.
Model Strict-F1 Ent-F1 RougeL-F Average
GPT4o 0.979 0.979 0.979 0.979
Claude3.5 0.967 0.978 0.970 0.972
DeepSeekV3 0.978 0.979 0.978 0.978
Llama3.1 0.978 0.984 0.978 0.980
Qwen2.5 0.980 0.988 0.980 0.982
Llama3.1-SLM 0.982 0.984 0.982 0.982
Qwen2.5-SLM 0.986 0.988 0.986 0.987
Overall Average 0.979 0.983 0.979 0.980
Table 15: Performance Consistency Scores for PII detection task across different metrics.
Model Task Naive Self-CoT Auto-CoT Self-Consistency PS-CoT Naive w/ Choice Average
GPT4o Query-Related 0.966 0.962 0.955 0.966 0.952 0.976 0.963
Masking 0.969 0.972 0.973 0.970 0.976 0.979 0.973
Llama3.1 Query-Related 0.972 0.969 0.967 0.960 0.966 0.972 0.968
Masking 0.976 0.977 0.973 0.970 0.966 0.967 0.972
Qwen2.5 Query-Related 0.976 0.965 0.974 0.969 0.972 0.980 0.973
Masking 0.971 0.972 0.972 0.967 0.969 0.973 0.971
Llama3.1-SLM Query-Related 0.954 0.956 0.953 0.948 0.951 0.967 0.955
Masking 0.963 0.969 0.968 0.965 0.964 0.962 0.965
Qwen2.5-SLM Query-Related 0.960 0.957 0.951 0.947 0.962 0.973 0.958
Masking 0.968 0.963 0.961 0.957 0.969 0.972 0.965
Overall Average 0.968 0.966 0.965 0.962 0.965 0.972 0.966
Table 16: Performance Consistency Scores for query-related detection and masking tasks by prompting strategy.
F.4.3 Representative Example from PII-Real
Figure 6 presents a representative instance from the PII-Real dataset, illustrating the comparable complexity and structure to synthetic samples. This example is derived from a real-world biographical profile. [^2] Original source: https://kimiyoung.github.io
User Description:
“Hi, I’m Zhilin Yang. I am working on a startup and I am the CEO of Moonshot AI. In 2019, I obtained my PhD degree from Carnegie Mellon University, advised by Ruslan Salakhutdinov and William W. Cohen. Prior to that, in 2015, I received my bachelor’s degree from Tsinghua University, advised by Jie Tang. I worked at Meta AI with Jason Weston, and Google Brain with Quoc V. Le.”
Query:
“In my current role, we’re early and juggling hiring with first product bets. What minimal weekly rhythm and decision rules keep us fast without creating chaos?”
PII Entities (Total: 16):
PER: Zhilin Yang, Ruslan Salakhutdinov, William W. Cohen, Jie Tang, Jason Weston, Quoc V. Le DEM: CEO, PhD degree, bachelor’s degree CODE: (none) ORG: Moonshot AI, Carnegie Mellon University, Tsinghua University, Meta AI, Google Brain DATETIME: 2019, 2015 LOC: (none) QUANTITY: (none) Query-Related PII: CEO, Moonshot AI
Figure 6: Example from PII-Real dataset showing a real-world biographical profile with comprehensive PII annotations. The query specifically targets career-related information, requiring identification of occupation and organizational affiliation while protecting other personal details.
F.5 Additional Results
Model Strict-F1 Ent-F1 RougeL-F
API-based Large Language Model
GPT4o 0.912 0.934 0.914
Claude3.5 0.887 0.911 0.889
DeepSeekV3 0.923 0.941 0.925
Open-source Large Language Model
Llama3.1 0.901 0.928 0.903
Qwen2.5 0.884 0.919 0.887
Open-source Small Language Model
Llama3.1-SLM 0.762 0.813 0.766
Qwen2.5-SLM 0.798 0.856 0.803
Table 17: Performance of baseline models on PII detection task (PII-Real dataset).
Method GPT4o Llama3.1 Qwen2.5 Llama3.1-SLM Qwen2.5-SLM
F1 RougeL-F F1 RougeL-F F1 RougeL-F F1 RougeL-F F1 RougeL-F
Basic Method
Naive 0.652 0.654 0.648 0.651 0.635 0.638 0.346 0.348 0.427 0.429
Advanced Method
Self-CoT 0.738 0.741 0.712 0.714 0.694 0.697 0.408 0.411 0.418 0.421
Auto-CoT 0.691 0.693 0.724 0.728 0.729 0.732 0.451 0.454 0.389 0.392
Self-Consistency 0.745 0.747 0.656 0.659 0.671 0.674 0.327 0.331 0.338 0.342
PS-CoT 0.683 0.685 0.673 0.676 0.689 0.691 0.368 0.372 0.468 0.471
w/ Extra Information
Naive w/ Choice 0.861 0.861 0.782 0.784 0.847 0.849 0.538 0.541 0.791 0.793
Table 18: Performance comparison on Query-Related PII Detection task (PII-Real dataset).
Method GPT4o Llama3.1 Qwen2.5 Llama3.1-SLM Qwen2.5-SLM
F1 RougeL-F F1 RougeL-F F1 RougeL-F F1 RougeL-F F1 RougeL-F
Basic Method
Naive 0.743 0.745 0.738 0.742 0.721 0.725 0.436 0.441 0.558 0.593
Advanced Method
Self-CoT 0.782 0.785 0.768 0.771 0.751 0.754 0.547 0.551 0.561 0.596
Auto-CoT 0.771 0.773 0.781 0.784 0.782 0.785 0.589 0.593 0.562 0.597
Self-Consistency 0.794 0.796 0.732 0.736 0.734 0.738 0.508 0.513 0.512 0.548
PS-CoT 0.758 0.761 0.745 0.748 0.753 0.756 0.498 0.503 0.578 0.614
w/ Extra Information
Naive w/ Choice 0.838 0.841 0.796 0.799 0.812 0.815 0.478 0.483 0.689 0.725
Table 19: Performance comparison on Query-Unrelated PII Masking task (PII-Real dataset).
GPT4o Llama3.1 Qwen2.5 Llama3.1-SLM Qwen2.5-SLM
Method F1 RougeL-F F1 RougeL-F F1 RougeL-F F1 RougeL-F F1 RougeL-F
Basic Method
Naive 0.600 0.602 0.611 0.614 0.596 0.603 0.240 0.333 0.405 0.413
Advanced Method
Self-CoT 0.675 0.681 0.638 0.643 0.626 0.632 0.354 0.362 0.392 0.397
PS-CoT 0.618 0.620 0.624 0.631 0.636 0.643 0.291 0.300 0.431 0.436
w/ Extra Information
F.6 Prompt Details
Appendix G PII Annotation System
G.1 PII Detection Interface
G.2 Query-Related PII Detection Interface
G.3 Query-Unrelated PII Masking Visualization
G.4 Annotation Guidelines
To ensure annotation consistency and quality, we established comprehensive guidelines for each task. Table 22 summarizes the core annotation instructions provided to annotators across the three tasks.
G.5 Quality Control and Inter-Annotator Agreement
To maintain high annotation quality, we implemented a rigorous quality control protocol. Each sample was independently annotated by multiple annotators, with disagreements resolved through majority voting or expert review. Regular review sessions were conducted to discuss challenging cases and update guidelines based on annotator feedback. For quality assurance, we randomly sampled 10% of the annotations for expert review.
Annotation Task Agreement (%) Fleiss’ $\kappa$
PII Detection 95.1 0.912
Query-Related Detection 91.5 0.873
Table 21: Inter-annotator agreement for PII detection and query-related detection tasks.
Table 21 presents the inter-annotator agreement results for the two primary annotation tasks, measured using both observed agreement rates and Fleiss’ kappa. We obtain kappa values of 0.912 for PII detection and 0.873 for query-related detection across all annotators.
Task Annotation Guidelines
PII Detection • Verify whether all PII entities in the user description are correctly identified.
• Click the Correct button if all entities are properly detected.
• Annotate missing PII entities by selecting text spans and assigning type, tag, and subject identifiers.
• Assign the next available letter (A, B, C) for entities belonging to new subject groups.
• Select minimal spans containing only the essential text representing each entity.
Query-Related PII Detection • Identify PII entities crucial for addressing the query from the provided options.
• Verify or correct the subject association (A, B, C, etc.) for each selected entity.
• Ensure that selected entities are both necessary and sufficient for query resolution.
• Use the review mechanism to validate your selection before submission.
Query-Unrelated PII Masking • Review the masked user description where query-unrelated PII entities are replaced with type tags.
• Verify that query-related entities are preserved in their original form.
• Confirm that masked entities are replaced with appropriate tags (e.g., <Nickname>, <Phone Number>).
• Evaluate whether the masked text maintains semantic coherence and readability.
• Assess the balance between privacy protection and query utility preservation.
Table 22: Comprehensive annotation guidelines for the three tasks in PII-Bench. Each task follows specific protocols to ensure consistency and quality across annotators.